chaylink05:

Anastasia uses the equation p = 0.7(rh + b) to estimate the amount of take-home pay, p, for h hours worked at a rate of r dollars per hour and any bonus received, b. What is an equivalent equation solved for h? h = (h equals StartFraction p Over 0.7 EndFraction minus b divided by r.– b)÷ r h = h equals StartFraction p Over 0.7 EndFraction minus b divided by r.– b ÷ r h = h equals left-parenthesis StartFraction p Over 0.7 EndFraction right-parenthesis divided by r minus b.÷ r – b h = h equals StartFraction p minus b Over 0.7 EndFraction divided by r. ÷ r

justus:

Welcome to Question Cove.

chaylink05:

yes thank you so much

justus:

You're most welcome :)

justus:

Good luck

Mercury:

the answer choice formatting is a little confusing but it's really just asking to solve the equation for h p = 0.7(rh + b) as a reminder, if you want to isolate a variable, you need to consider what's being done to the variable, and do the opposite for example, since the entire expression on the right side is being multiplied by 0.7, you can divide by 0.7 on both sides to get p/0.7 = rh + b and so on. the next step would be to eliminate the +b (how might you do that?) and isolate h from rh (how might you do that?